A restaurant has some pancake batter that is too thin. It weighs 12 kilograms and contains 85% flour. How many kilograms of flour should be added so that the final mixture is 90% flour?
Answer by gonzo(654)

there are at least a couple of ways of looking at it.
here's one.
-----------------------------------
F = .85M (F = flour)
W = .15M (W = water)
M = 12 (M = mixture)
F = 10.2 kg
W = 1.8 kg
since the amount of flour needs to be raised, we can keep the amount of water constant at 1.8 kg.
1.8 kg of water in the new mixture represents 10% of the mixture (90% flour).
1.8 kg of water / .10 = 18 kg of mixture
new mixture must be 18 kg if 1.8 kg of water is 10% of it.
amount of flour in the new mixture becomes 18 - 1.8 = 16.2 kg
16.2 - 10.2 kg of flour needs to be added to the mixture = 6 kg of flour.
answer is 6 kg of flour needs to be added
-------------------------------------
here's the other.
-------------------------------------
if .85 of the mixture is flour and .15 of the mixture is water, then the ratio of flour to water is represented by (85/15)
to test you take 85/15 * 1.8 kg water = 10.2 kg of flour - checks out ok.
since the new mixture must be 90% flour and 10% water, the ratio of flour to water in the new mixture must be (9/1).
since the amount of water will remain the same, (9/1)*1.8 = 16.2 kg of flour required in the new mixture.
16.2 - 10.2 = 6 kg.
--------------------------------------